OpenStudy (anonymous):

The absoption rate of medication is an application of a special type of series, the infinite geometric series. an=3(1/2)^(n-1) First term of the sequence is 3, and the common ratio is 1/2. Find the sum of the first 3 terms, or S3. Either add the first three terms, or use the formula Sn=a1-a1^r^n/1-r. PLEASE HELP?!

OpenStudy (anonymous):

it is given that a1 = 3 a2 = 3(1/2)^(2-1) = 3/2 a3= 3/4 add the first three terms 3+3/2+3/4 = 3(1+1/2+1/4) = 21/4

OpenStudy (anonymous):

S3 = 3 (1- 1/2^3) / 1 - 1/2 = 3 * 7/8 / 1/2 = 42/8 = 21/4 6 6